OpenStudy (anonymous):

phase shift? y=-2cos(4x+pi)

OpenStudy (anonymous):

pi/4?

OpenStudy (anonymous):

great jims, are you going to help?

OpenStudy (mathmath333):

i think its \(\Large \bf \dfrac{\pi}{4}\) not sure though

OpenStudy (anonymous):

how did u get it?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{y=-2cos(4x+pi)\\ y=-2cos(4\color{red}{(x+\dfrac{\pi}{4})})}\end{align}\)